# Structural Nested Mean Models for Modified Treatment Policies

### v1

@@ 1 Introduction
The structure of the paper is as follows. In Section 2, we summarize notation and review definitions of MTPs and SNMMs. Once defined, we also provide some hypothetical examples of when SNMMs for MTP effects might be of interest. In Section 3, we provide identification results and Neyman orthogonal estimators under exchangeability assumptions. In Section 4, we provide identification results under parallel trends assumptions. In Section 5, we provide simulations showing that our estimators are unbiased and that the sandwich variance estimates provide nominal coverage. In Section 6, we present a real data analysis estimating the effects of shifting mobility from its natural level on subsequent county level Covid-19 incidence. In Section 7, we conclude and discuss some directions for future work.
@@ 3.1 Point Exposure Setting
Algorithm 1 computes the estimator from Theorem 1.
@@ 3.2 Time varying treatments
Assume Sequential consistency, positivity, and MTP exchangeability hold (i.e. the assumptions of Section 5 of Richardson and Robins [2013]). Consider the score
@@
    Asymptotic variance Suppose $m=d$ and $\psi$ is estimated by the root $\hat{\psi}$ of the cross-fitted sample moment $n^{-1}\sum_{i=1}^{n}\hat{\phi}_{i}(\psi)=0$, with nuisances learned on held-out folds and consistent in $L_{2}$. Let
    $G:=E\!\left[\frac{\partial}{\partial\psi^{\top}}\phi(O;\psi^{\ast},\eta^{\ast})\right]\in R^{d\times d},\qquad\Sigma:=Var\!\big(\phi(O;\psi^{\ast},\eta^{\ast})\big)\in R^{d\times d}.$
@@
4 Identification and Parametric Estimation Under Parallel Trends Assumptions
@@ 4.1 Point Exposure Setting
where $\mu^{d}(L,A;\beta)$ is a model for $\mu^{d}(L,A)$ smooth in an $r$-dimensional parameter $\beta$ (equal to $\beta^{*}$ under the true law) and $q(A,L)$ and $b(A,L)$ are conformable analyst selected index functions. Derivation of Neyman orthogonal estimators is left for future work.
@@ 4.2 Time-Varying Treatments
where $\mu_{mk}^{d}(\bar{A}_{m},\bar{L}_{m};\beta)$ are models for $\mu_{mk}^{d}(\bar{A}_{m},\bar{L}_{m})$ smooth in an $r$-dimensional parameter $\beta$ (equal to $\beta^{*}$ under the true law) and $q_{mk}(\bar{A}_{m},\bar{L}_{m})$ and $b_{mk}(\bar{A}_{m},\bar{L}_{m})$ are conformable analyst selected index functions. Derivation of Neyman orthogonal estimators is again left for future work.
@@ 7 Discussion and Future Work
Furthermore, economists who commonly apply DiD can now study MTPs (albeit only using parametric nuisance models for the time being) under similar parallel trends assumptions. For example, they can estimate the effect of increasing minimum wage by 1 dollar more than it actually increased, and how this effect varies with the amount it actually increased. This is not an estimand that can be targeted by standard DiD.
There are many directions for future work. First, it remains to derive Neyman orthogonal estimators under parallel trends assumptions. Given past modifications of SNMM results to accommodate parallel trends [Shahn et al., 2022], we believe this should be achievable. Second, SNMMs can also be estimated under instrumental variable assumptions [Robins, 1994b], and instrumental variable estimation of MTPs would also be of interest. Third, SNMM variants have been developed to estimate effects on survival [Picciotto et al., 2012] and binary [Wang et al., 2023] outcomes. Effects of MTPs on survival and binary outcomes are of course of interest as well. Fourth, sensitivity analysis for SNMMs [Robins, 2004, Robins et al., 2000] is well developed and would hopefully easily port over to the MTP setting, but that should be confirmed. Finally, marginal SNMMs that model heterogeneity as a function of a subset of covariates required for adjustment have been developed elsewhere, and it might be useful to transport them to the MTP setting, too.

### v2

@@ 1 Introduction
The structure of the paper is as follows. In Section 2, we summarize notation and review definitions of MTPs and SNMMs. Once defined, we also provide some hypothetical examples of when SNMMs for MTP effects might be of interest. In Section 3, we provide identification results and Neyman orthogonal estimators under exchangeability assumptions. In Section 4, we do the same under parallel trends assumptions. In Section 5, we provide simulations showing that our estimators are unbiased and that the sandwich variance estimates provide nominal coverage. In Section 6, we present a real data analysis estimating the effects of shifting mobility from its natural level on subsequent county level Covid-19 incidence. In Section 7, we conclude and discuss some directions for future work.
@@ 3.1 Point Exposure Setting
Algorithm 1 computes the estimator from Theorem 1. Note that the theorem requires only that $T_{g}^{\dagger}$ satisfies the density-weighted adjoint identity (10). The familiar formulas for the settings where $g$ is a continuous bijection or discrete are convenient instances but not required for the theorem to hold.
@@ 3.2 Time varying treatments
Assume Sequential consistency, positivity, and MTP exchangeability (i.e. the assumptions of Section 5 of Richardson and Robins [2013]) hold. Consider the score
@@
    Asymptotic variance Suppose $\psi$ is estimated by the root $\hat{\psi}$ of the cross-fitted sample moment $n^{-1}\sum_{i=1}^{n}\hat{\phi}_{i}(\psi)=0$, with nuisances learned on held-out folds and consistent in $L_{2}$. Let
    $G:=E\!\left[\frac{\partial}{\partial\psi^{\top}}\phi(O;\psi^{\ast},\eta^{\ast})\right],\qquad\Sigma:=Var\!\big(\phi(O;\psi^{\ast},\eta^{\ast})\big).$
@@
4 Identification and Estimation Under Parallel Trends Assumptions
@@ 4.1 Point Exposure Setting
where $\mu^{d}(L,A;\beta)$ is a model for $\mu^{d}(L,A)$ smooth in an $r$-dimensional parameter $\beta$ (equal to $\beta^{*}$ under the true law) and $q(A,L)$ and $b(A,L)$ are conformable analyst selected index functions.
@@ 4.2 Time-Varying Treatments
where $\mu_{mk}^{d}(\bar{A}_{m},\bar{L}_{m};\beta)$ are models for $\mu_{mk}^{d}(\bar{A}_{m},\bar{L}_{m})$ smooth in an $r$-dimensional parameter $\beta$ (equal to $\beta^{*}$ under the true law) and $q_{mk}(\bar{A}_{m},\bar{L}_{m})$ and $b_{mk}(\bar{A}_{m},\bar{L}_{m})$ are conformable analyst selected index functions.
4.3 Neyman orthogonal estimators
To derive Neyman orthogonal estimators enabling machine learning estimation of nuisance functions, we proceed completely analogously to the (sequential) exchangeability case. In fact, we can simply substitute the difference $\Delta Y_{t}\equiv Y_{t}-Y_{t-1}$ for $Y_{t}$ and the regressions $\mu_{mk}^{d}(H_{m},A_{m})=E[\Delta Y_{k}|H_{m},A_{m}]$ for $\mu_{mk}(H_{m},A_{m})$ in the scores from Theorems 1 and 2. The proofs of those theorems then go through exactly as before. Here, we state the theorem for the general longitudinal parallel trends case.
Theorem 3 (Neyman–orthogonal estimator under parallel trends, longitudinal).
Set
$\mu_{t}^{d}(h_{t},a_{t}):=\mathbb{E}[\Delta Y_{t}\mid H_{t}=h_{t},A_{t}=a_{t}].$
For each $t$, equip functions with $\langle u,v\rangle_{t}:=\mathbb{E}\!\big[\int u(H_{t},a_{t})v(H_{t},a_{t})f_{A_{t}\mid H_{t}}(a_{t}\mid H_{t})\,da_{t}\big].$ Let $T_{g_{t}}$ act by $(T_{g_{t}}q_{t})(h_{t},a_{t})=q_{t}(h_{t},g_{t}(h_{t},a_{t}))$ with adjoint $T_{g_{t}}^{\dagger}$ under $\langle\cdot,\cdot\rangle_{t}$. Choose $q_{t}(H_{t},A_{t}):=s_{t}(H_{t})$ and set $\tilde{q}_{t}:=T_{g_{t}}^{\dagger}q_{t}$. Let $\Gamma_{t}(H_{t};\psi_{t})$ be linear in $s_{t}(H_{t})$.
Assume for each $t$: $\mathbb{E}[\Delta Y_{t}(0)\mid H_{t},A_{t}]=\mathbb{E}[\Delta Y_{t}(0)\mid H_{t}]$. Further assume sequential consistency, positivity, and regularity conditions.
Define the time-$t$ score
$\phi_{t}=(q_{t}-\tilde{q}_{t})\{\Delta Y_{t}-\mu_{t}^{d}(H_{t},A_{t})\}+s_{t}(H_{t})\{\mu_{t}^{d}(H_{t},A_{t})-\mu_{t}^{d}(H_{t},g_{t}(H_{t},A_{t}))+\Gamma_{t}(H_{t};\psi_{t})\}.$
Let $\phi=(\phi_{0}^{\top},\dots,\phi_{T-1}^{\top})^{\top}$ and $\psi=(\psi_{0}^{\top},\dots,\psi_{T-1}^{\top})^{\top}$.
Then:
  1. (i)
    Identification. For each $t$, there is a unique $\psi_{t}^{*}$ s.t. $\mathbb{E}[s_{t}\{\mu_{t}^{d}(H_{t},A_{t})-\mu_{t}^{d}(H_{t},g_{t}(H_{t},A_{t}))+\Gamma_{t}(H_{t};\psi_{t})\}]=0$, and $\mathbb{E}[\phi_{t}(\psi_{t}^{*},\eta_{t}^{*})]=0$. Hence $\mathbb{E}[\phi(\psi^{*},\eta^{*})]=0$.
  2. (ii)
    Neyman orthogonality. For any regular paths $\eta_{t,\varepsilon}=(\mu_{t,\varepsilon}^{d},f_{t,\varepsilon})$ through $\eta_{t}^{*}$, $\left.\frac{d}{d\varepsilon}\mathbb{E}[\phi_{t}(\psi_{t}^{*},\eta_{t,\varepsilon})]\right|_{\varepsilon=0}=0$ for each $t$.
  3. (iii)
    Asymptotic normality. Suppose $\psi$ is estimated by the root $\hat{\psi}$ of the cross-fitted sample moment $n^{-1}\sum_{i=1}^{n}\hat{\phi}_{i}(\psi)=0$, with nuisances learned on held-out folds and consistent in $L_{2}$. Let
    $G:=E\!\left[\frac{\partial}{\partial\psi^{\top}}\phi(O;\psi^{\ast},\eta^{\ast})\right],\qquad\Sigma:=Var\!\big(\phi(O;\psi^{\ast},\eta^{\ast})\big).$
    Then under standard regularity conditions,
    and
    $V\;=\;G^{-1}\,\Sigma\,(G^{-1})^{\top}.$
@@ 7 Discussion and Future Work
Furthermore, economists who commonly apply DiD can now study MTPs under similar parallel trends assumptions. For example, they can estimate the effect of increasing minimum wage by 1 dollar more than it actually increased, and how this effect varies with the amount it actually increased. This is not an estimand that can be targeted by standard DiD.
There are many directions for future work. First, SNMMs can be estimated under instrumental variable assumptions [Robins, 1994b], and instrumental variable estimation of MTPs would also be of interest. Second, SNMM variants have been developed to estimate effects on survival [Picciotto et al., 2012] and binary [Wang et al., 2023] outcomes. Effects of MTPs on survival and binary outcomes are of course of interest as well. Third, sensitivity analysis for SNMMs [Robins, 2004, Robins et al., 2000] is well developed and would hopefully easily port over to the MTP setting, but that should be confirmed. Finally, marginal SNMMs that model heterogeneity as a function of a subset of covariates required for adjustment have been developed elsewhere, and it might be useful to transport them to the MTP setting, too.